OpenStudy (anonymous):

find the equation of straight line joining the point of intersection x^2+y^2=1 and y2=4x is given by

OpenStudy (anonymous):

is that 2y or y^2 = 4x?

hartnn (hartnn):

solve them simultaneously plug in y^2 =4x in x^2+y^2 =1 you'll get a quadratic in 'x'

OpenStudy (anonymous):

its y^2=4x @williamnz

OpenStudy (anonymous):

@hartnn but the options of answer are x=2 x=-2 y=-2 y=2

OpenStudy (anonymous):

i have done in the way u told me so @hartnn the answer is x=2 right?

hartnn (hartnn):

thats because you still need to solve that quadratic! x^2+4x =1 how did you get x=2 ?? can you show me your steps ?

hartnn (hartnn):

can you please verify your question ? it does not seem the correct answer is in the options

OpenStudy (anonymous):

i put the value of y^2+4x in equation 1 then apply the quadratic formula and i get two answers x=2 and x=-6

OpenStudy (anonymous):

yeh it is

hartnn (hartnn):

this ? \(x^2+y^2 =1 \\ y^2 =4x\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y^2=4x x^2+y^2=12 then x^2+4x-12+0 applying quadratic equation i get two answer i ignore the minus value and put my answer as x=2

hartnn (hartnn):

oh it is 12 and not 1

hartnn (hartnn):

then yes x=2 is correct :)

OpenStudy (anonymous):

thanku @hartnn

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

which website is best for preparing apptitude test?? @hartnn

OpenStudy (anonymous):

which of the following line is parallel to line 3x-12y=12 x-4y=-1 x+4y=-1 -x-4y=-1 3x+12y=0

hartnn (hartnn):

find the slope of 3x-12y =12

OpenStudy (anonymous):

then? the slope is 1/4

hartnn (hartnn):

right now "Parallel lines have same slope!" so check which line from your choices have slope =1/4

OpenStudy (anonymous):

ok according to you which option is correct?

hartnn (hartnn):

can you find the slope of x-4y=-1 ?

OpenStudy (anonymous):

yeh its 1/4 :)

hartnn (hartnn):

then thats your answer :)

OpenStudy (anonymous):

hmmm thanks;)

hartnn (hartnn):

welcome ;)

OpenStudy (anonymous):

y varies directly as cube of x. when x=0.5, y=2.5. if y=540 then x=

hartnn (hartnn):

cube of x : \(x^3 \) y varies directly as cube of x \(y = kx^3 \) from x= 0.5, y= 2.5 find the value of k

OpenStudy (anonymous):

k=20x^3 and the answer must be x=3?

hartnn (hartnn):

y = 20x^3 and yes x=3 is correct :)

OpenStudy (anonymous):

mean i did it rightly but then i thought my answer is wrong so i post it:P

hartnn (hartnn):

its always good to verify :)

OpenStudy (anonymous):

hmm:) gonna ask some more questions

hartnn (hartnn):

sure, go ahead :) just a tip : make a new post for each question!

OpenStudy (anonymous):

oh ok:)